OpenStudy (anonymous):

Write an equation for each translation of y=|x|. 6.5 units up y + 6.5 = | x | y = | 6.5x | y = | x | + 6.5 y = | x | – 6.5

OpenStudy (anonymous):

jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

we just tack on a `+6.5` to the end to shift the graph up 6.5 units

jimthompson5910 (jim_thompson5910):

so it would be `y = | x | + 6.5`

OpenStudy (anonymous):

I have another one

jimthompson5910 (jim_thompson5910):

hint: point slope form \[\LARGE y-y_{1} = m(x-x_{1})\]

OpenStudy (anonymous):

C

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

look at point slope form carefully

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

B

jimthompson5910 (jim_thompson5910):

sounds like you're randomly guessing at this point

OpenStudy (anonymous):

let me think about it.

OpenStudy (anonymous):

y=-6 so D?

jimthompson5910 (jim_thompson5910):

yeah it should be y-(-6) which turns into y+6

OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (–3, –7); m = -6/5 y – 7 = -6/5(x – 3) y + 7 = -6/5(x + 3) y + 3 = -6/5(x + 7) y – 7 = -6/5(x + 3)

jimthompson5910 (jim_thompson5910):

what are your thoughts on this one?

OpenStudy (anonymous):

If it like the other ones it would be C

jimthompson5910 (jim_thompson5910):

you mixed up x and y though

OpenStudy (anonymous):

so I flip the 7 and 3?

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (anonymous):

B?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Thanks for the help